OpenStudy (anonymous):

If (2m +7n)/ (4m^2-49n^2) = 7, what is the value of 2m - 7n?

OpenStudy (callisto):

\[\frac{2m +7n}{4m^2-49n^2} = 7\]\[\frac{2m +7n}{(2m)^2-(7n)^2} = 7\]Use difference of two squares to factor it. \(a^2-b^2=(a-b)(a+b)\)\[\frac{2m +7n}{(2m-7n)(2m+7n)} = 7\]Cancel the common factor in the fraction, what do you get so far?

OpenStudy (anonymous):

1/(2m-7n) = 7

OpenStudy (callisto):

Multiply both sides by 2m-7n, what do you get?

OpenStudy (anonymous):

14m-49n = 1 so 2m- 7n = 1/7 thanks :)

OpenStudy (callisto):

Welcome :D